OpenStudy (anonymous):

A spaceship moving with an initial velocity of 58.0 meters/second experiences a uniform acceleration and attains a final velocity of 153 meters/second. What distance has the spaceship covered after 12.0 seconds?

OpenStudy (anonymous):

7.9 or 8?\[(vf-vi) \div t=d\]

OpenStudy (anonymous):

(153 m/s - 58 m/s) / 12s = 95m/s / 12s = 7.9

OpenStudy (anonymous):

7.9 meters

OpenStudy (irishboy123):

use \(v = u + at\) to get the acceleration then \(v^2 = u^2 + 2ax\) to get the distance travelled

OpenStudy (anonymous):

or\[v_f^2 = v_i^2 + 2a \Delta d\]